# Penrose’s eight-conic theorem via Penrose’s eight-quadric theorem

### v3

@@ 1 Introduction
Consider a pair of lines in the plane (call them $\ell$ and $\ell^{\prime}$). Assign three arbitrary points on each line ($A,B,C$ on $\ell$ and $A^{\prime},B^{\prime},C^{\prime}$ on $\ell^{\prime}$). Join these three pairs of points in a crossing manner, giving us three meeting points (namely the meet of lines $AB^{\prime}$ and $A^{\prime}B$, the meet of lines $BC^{\prime}$ and $B^{\prime}C$, and the meet of lines $CA^{\prime}$ and $C^{\prime}A$). Then these three points must be collinear.
Consider a pair of points in the plane (call them $L$ and $L^{\prime}$). Draw three lines through each point ($a,b,c$ through $L$ and $a^{\prime},b^{\prime},c^{\prime}$ through $L^{\prime}$). Meet these three pairs of lines in a diagonal manner, which are joined into three lines (namely the join of points $ab^{\prime}$ and $a^{\prime}b$, the join of points $bc^{\prime}$ and $b^{\prime}c$, and the join of points $ca^{\prime}$ and $c^{\prime}a$). Then these three lines must concur.
@@
Figure 1: Pappus’ theorem (340 AD) (left) and its dual theorem (right).
A fascinating quality about Pappus’ theorem is that the statement of the theorem requires very few structures about the plane, that two distinct points can be joined into a unique line, and two distinct lines meet at a unique point (possibly at the line at infinity), like those rules in perspective drawings. No lengths, angles, or parallelisms are involved. This is a projective geometric theorem. Note that one may swap “points” and ”lines” and swap “meet” and “join” in a planar projective geometric theorem to obtain the dual theorem, as shown above.
It is fair to say that neither Pappus’ theorem and its dual are obvious. A natural question is whether there is a general pattern to theorems like Pappus’, so that upon this understanding, Pappus’ theorem becomes more intuitive and easier to discover.
Blaise Pascal (1623–1662), at age 16, discovered a remarkable generalization of Pappus’ theorem, known as Pascal’s mystic hexagram [DC20]. The initial line pair of Pappus’ theorem, as well as the initial point pair of the dual Pappus theorem, is a degeneration of a conic section. After the line pair is replaced by an arbitrary conic, the rest of the theorem statement remains true (Figure 2).
@@
Consider a conic in the plane. Assign six points on the conic (labeled $A,B,C,A^{\prime},B^{\prime},C^{\prime}$). Join these three pairs of points in a crossing manner, giving us three meeting points (namely the meet of lines $AB^{\prime}$ and $A^{\prime}B$, the meet of lines $BC^{\prime}$ and $B^{\prime}C$, and the meet of lines $CA^{\prime}$ and $C^{\prime}A$). Then these three points must be collinear.
Consider a conic in the plane. Draw six tangent lines of the conic (labeled $a,b,c,a^{\prime},b^{\prime},c^{\prime}$). Meet these three pairs of lines in a diagonal manner, which are joined into three lines (namely the join of points $ab^{\prime}$ and $a^{\prime}b$, the join of points $bc^{\prime}$ and $b^{\prime}c$, and the join of points $ca^{\prime}$ and $c^{\prime}a$). Then these three lines must concur.
@@
In 1848, George Salmon published a textbook on conic sections, which include a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal theorem are degeneration of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs by three conics in double contact with a fourth, the remainder of Pappus’ theorem statement still holds (Figure 3).
@@
If three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ have each in double contact with a fourth $\mathbf{{S}}_{0}$, then the three intersection points, each from a pair of common tangents between $\mathbf{{T}}^{1}\mathbf{{T}}^{2}$, between $\mathbf{{T}}^{2}\mathbf{{T}}^{3}$, and between $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$, are collinear.
If three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ have each in double contact with a fourth $\mathbf{{S}}_{0}$, then a common chord between $\mathbf{{T}}^{1}\mathbf{{T}}^{2}$, a common chord between $\mathbf{{T}}^{2}\mathbf{{T}}^{3}$, and a common chord between $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$, are concurrent. [^6] The original statement is “If three conics have each in double contact with a fourth, then six of their chords of intersection form a complete quadrilateral.”
The reader may have observed a pattern of this progression of generalizing Pappus’ theorem. A line pair and a point pair are replaced by conic sections. The incidence relations of point pairs on a conic or line pairs tangent to a conic are both unified by two conics in double contact.
@@
Figure 4: The main theorem—the eight conic theorem—first enunciated by Roger Penrose in around 1950. The right figure illustrates an instance of the configuration where all 8 conics are ellipses, shown next to the corresponding cube graph of conics and double contact relationships.
@@
This article shows a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last author of this article, in around 1950 during his college years. We refer to it Penrose’s eight-conic theorem (Figure 4):
@@
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, then (upon a mild consistency condition on the cube face) there exists an eighth conic which completes the cube.
@@
Namely, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each in double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are respectively in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$ (chosen in a consistent branch of a double contact family), then there exists a conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$.
@@
Here, the consistency condition of the cube face is that the four chords of contact from the four conics have a common point. This condition is automatically guaranteed if one picks a consistent branch of double contact conics with two given conic (out of 3 branches).
@@
This article provides a completion of one of Roger Penrose’s proof ideas for the eight-conic theorem. This idea emerged during his graduate studies at Cambridge, which is that the eight-conic configuration is a planar slice of eight quadrics in ring contact in 3D. (Two quadrics have ring contact if they touch along a conic.) It turns out that the following Penrose’s eight-quadric theorem is much more intuitive to see how it has to hold, as illustrated later in this article:
@@
If quadrics in 3D are assigned to seven of the vertices of a combinatorial cube such that quadrics connected by an edge are in ring contact, then (upon a mild consistency condition on the cube face) there exists an eighth quadric which completes the cube.
@@
The face consistency condition is that the four planes containing the four rings of the ring contacts must meet along a common axis. Unlike in 2D, there are generally no incompatible quadrics in 3D when completing the vertices of the cube, except in very special situations.
@@
The article is structured as follows. Section 2 provides the necessary background on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.3, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 4. Section 5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. The final sections of the article provide references to related works and future outlooks.
@@ 2.1 Conics and the projective plane
We recall that a point $P$ in the projective plane $\mathbb{R}P^{2}$ can be represented by an equivalence class of vectors in a 3 dimensional real [^7] We work primarily over the real field but allow complex coefficients when they arise through algebraic closure. vector space $\mathtt{V}$:
@@
for $v\in\mathtt{V}^{*}\setminus\{0\}$. We could equally well have started by describing lines according to equation (3) and proceeded to points by describing them as the sets of incident lines:
@@
thus the roles of the two are completely interchangeable: the principle of duality. The symmetrical relationship of the two can be expressed with the notation
@@
A point-line correlation is an invertible map $\mathbb{R}P^{2}\mapsto{\mathbb{R}P^{2}}^{*}$ induced by a linear transformation $\mathtt{V}\mapsto\mathtt{V}^{*}$. The fundamental theorem of projective geometry states that this is the most general type of mapping between the two which always sends any three collinear points to three copunctual lines. A line-point correlation is a similarly defined map ${\mathbb{R}P^{2}}^{*}\mapsto{\mathbb{R}P^{2}}$.
@@
The linear transformation defining a point-line correlation can then be represented by a matrix $A$ relative to the bases $\mathcal{B}$ and $\mathcal{B}^{*}$, and vice versa for a line-point correlation. The matrix $cA$ for $c\neq 0$ clearly represents the same correlation (note however that changing the relative scaling of the vectors representing $P_{j}$ leads to a different matrix). A point-line correlation has two naturally associated line-point correlations: its inverse which has matrix $A^{-1}$ and the mapping of a line $p$ to the intersection point of any two lines mapped to by two points lying on $p$ which is well defined and has matrix $\mathrm{cof}A$, the cofactor matrix of $A$. By Cramer’s rule,
@@
where the superscript T denotes the transpose, therefore these two associated line-point correlations coincide exactly when the point-line correlation is represented by a symmetric matrix $A=A^{T}$ which is known as a polarity.
If $A$ is the matrix of a polarity, and $P$ is a point, which we identify with the column vector of a representative in the basis $(P_{0},P_{1},P_{2})$, then the line $p$ polar to $P$ is represented by a column vector, which we again identify with $p$ itself, in the basis $(p_{0},p_{1},p_{2})$:
$p=AP.$
@@
$p^{T}P=P^{T}AP=0.$
The locus of such points is a regular point-wise conic which is uniquely defined by the polarity and vice versa, both having a unique representative symmetric matrix (up to scaling) in a given basis. The line to which a point of a point-wise conic is both polar and incident is its tangent line. Dual to this, the locus of lines incident to their polar points is called a regular linewise conic and we refer to these points as the turning points of the linewise conic. When convenient and no ambiguity can result, we sometimes identify the matrix representing a conic in a given basis with the conic itself.
Given a regular point-wise conic, there is a uniquely associated regular linewise conic, the lines of the latter being the tangent lines of the points of the former. If the first conic is represented by the matrix $A$, the second one is represented by the inverse matrix $A^{-1}$. A degenerate point-wise conic is a conic defined in the same way but by a non-invertible symmetric matrix. In this case, three collinear points will still map to three copunctual lines when all images are defined, however, for any point $P$ whose representatives lie in the kernel of the defining matrix $A$, the image of $P$ under the degenerate polarity is geometrically undefined, since the zero vector does not represent any line in ${\mathbb{R}P^{2}}^{*}$, moreover, if $Q$ is any other point, then
@@
  • If $\mathrm{ker}A$ has dimension $1$ then it constitutes one equivalence class in $\mathbb{R}P^{2}$ so there is a unique such point. In this case the points of a degenerate point-wise conics constitute a line pair with $P$ as their meeting point. If these two lines are called $p$ and $q$ then the matrix is given by
@@
  • If $\mathrm{ker}A$ has dimension $2$ then it constitutes the representatives of the points on a line in the projective plane. In this case the points of a degenerate point-wise conics constitute a line, referred to for algebraic reasons as a doubled line. If this line is called $p$ then the matrix is given by
@@
Dual to this, degenerate linewise conics can be defined:
  • If $\mathrm{ker}A$ has dimension $1$ then it constitutes one equivalence class in ${\mathbb{R}P^{2}}^{*}$ so there is a unique line $p$ represented by vectors in the kernel. In this case the lines of a degenerate linewise conics constitute a point pair with $p$ as their joining line. If these to points are called $P$ and $Q$ then the matrix is given by
@@
  • If $\mathrm{ker}A$ has dimension $2$ then it constitutes the representatives of the lines through a point in the projective plane. In this case the lines of a degenerate linewise conics constitute a point, referred to for algebraic reasons as a doubled point. If this point is called $P$ then the matrix is given by

### v4

@@ 1 Introduction
Consider a pair of lines in the plane (call them $\ell$ and $\ell^{\prime}$). Assign three arbitrary points on each line ($A,B,C$ on $\ell$ and $A^{\prime},B^{\prime},C^{\prime}$ on $\ell^{\prime}$). Join these three pairs of points in a crossing pattern, giving three meeting points (namely, the meet of lines $AB^{\prime}$ and $A^{\prime}B$, the meet of lines $BC^{\prime}$ and $B^{\prime}C$, and the meet of lines $CA^{\prime}$ and $C^{\prime}A$). Then these three points lie on a line.
Consider a pair of points in the plane (call them $L$ and $L^{\prime}$). Draw three lines through each point ($a,b,c$ through $L$ and $a^{\prime},b^{\prime},c^{\prime}$ through $L^{\prime}$). Meet these three pairs of lines in a diagonal pattern, giving three lines (namely the join of points $ab^{\prime}$ and $a^{\prime}b$, the join of points $bc^{\prime}$ and $b^{\prime}c$, and the join of points $ca^{\prime}$ and $c^{\prime}a$). Then these three lines pass through a point.
@@
Figure 1: Pappus’s theorem (340 AD) (left) and its dual theorem (right).
A fascinating quality of Pappus’s theorem is that the statement of the theorem requires very little geometric structure, simply that two distinct points can be joined with a unique line, and two distinct lines meet in a unique point (possibly at the line at infinity), reminiscent of the practice of perspective drawing. No lengths, angles, or parallelisms are involved. This is a theorem of projective geometry. Note that one may swap “points” and ”lines” and swap “meet” and “join” in a planar projective geometric theorem to obtain the dual theorem, as shown above.
It is fair to say that neither Pappus’s theorem nor its dual are obvious. A natural question is whether there is a general pattern to theorems like Pappus’s, that could shed light on Pappus’s theorem, making it more intuitive and easier to discover.
Blaise Pascal (1623–1662), at age 16, discovered a remarkable generalization of Pappus’s theorem, known as Pascal’s mystic hexagram [DC20]. The initial line pair of Pappus’s theorem, as well as the initial point pair of the dual Pappus theorem, is a degeneration of a conic section. After replacing the pair of lines by an arbitrary conic, the rest of the statement of the theorem remains true (Figure 2).
@@
Consider a conic in the plane. Choose six points on the conic (labeled $A,B,C,A^{\prime},B^{\prime},C^{\prime}$). Join these three pairs of points in a crossing pattern, giving three meeting points (namely, the meet of lines $AB^{\prime}$ and $A^{\prime}B$, the meet of lines $BC^{\prime}$ and $B^{\prime}C$, and the meet of lines $CA^{\prime}$ and $C^{\prime}A$). Then these three points lie on a line.
Consider a conic in the plane. Draw six lines tangent to the conic (labeled $a,b,c,a^{\prime},b^{\prime},c^{\prime}$). Meet these three pairs of lines in a diagonal manner, giving three joining lines (namely the join of points $ab^{\prime}$ and $a^{\prime}b$, the join of points $bc^{\prime}$ and $b^{\prime}c$, and the join of points $ca^{\prime}$ and $c^{\prime}a$). Then these three lines go through a point.
@@
In 1848, George Salmon published a textbook on conic sections, which includes a further generalization of the Pascal–Brianchon theorem [Sal17, §264], rediscovered by Evelyn, Money-Coutts, and Tyrrell in 1974 [EMCT74, §2.3]. The three point pairs $AA^{\prime}$, $BB^{\prime}$, $CC^{\prime}$ of Pascal’s theorem are degenerations of thin ellipses that are in double contact with the initial conic. And the lines $AB^{\prime}$ and $A^{\prime}B$ are the limit of a pair of common tangents between the two thin ellipses collapsing to $AA^{\prime}$ and $BB^{\prime}$. After replacing the three point pairs by three conics in double contact with a fourth, the remainder of Pappus’s theorem statement still holds (Figure 3).
@@
If three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ are each in double contact with a fourth $\mathbf{{S}}_{0}$, then the three intersection points of pairs of common tangents (namely, between $\mathbf{{T}}^{1}\mathbf{{T}}^{2}$, between $\mathbf{{T}}^{2}\mathbf{{T}}^{3}$, and between $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$), lie on a line.
If three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ are each in double contact with a fourth $\mathbf{{S}}_{0}$, then a common chord between $\mathbf{{T}}^{1}\mathbf{{T}}^{2}$, a common chord between $\mathbf{{T}}^{2}\mathbf{{T}}^{3}$, and a common chord between $\mathbf{{T}}^{3}\mathbf{{T}}^{1}$, all pass through a point. [^6] The original statement is “If three conics are each in double contact with a fourth, then six of their chords of intersection form a complete quadrilateral.”
The reader may have observed a pattern of this progression of generalizing Pappus’s theorem. Line pairs and point pairs are replaced by conic sections. The incidence relations of point pairs on a conic or line pairs tangent to a conic are both generalized by two conics in double contact.
@@
Figure 4: The main theorem—the eight-conic theorem—first enunciated by Roger Penrose in around 1950. The right figure illustrates an instance of the configuration where all 8 conics are ellipses, shown next to the corresponding cube graph of conics and double contact relationships.
@@
This article shows a “final” generalization of this progression of projective geometric theorems, first stated by Roger Penrose, the last author of this article, in around 1950 during his undergraduate years. We refer to it Penrose’s eight-conic theorem (Figure 4):
@@
If conics are assigned to seven of the vertices of a combinatorial cube such that conics connected by an edge are in double contact, then (subject to a mild consistency condition on the cube faces) there exists an eighth conic which completes the cube.
@@
That is, if three conics $\mathbf{{T}}^{1},\mathbf{{T}}^{2},\mathbf{{T}}^{3}$ each in double contact with a fourth $\mathbf{{S}}_{0}$, and another three conics $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$ are, respectively, in double contact with $\mathbf{{T}}^{2},\mathbf{{T}}^{3}$, with $\mathbf{{T}}^{3},\mathbf{{T}}^{1}$, and with $\mathbf{{T}}^{1},\mathbf{{T}}^{2}$ (chosen in a consistent branch of a double contact family), then there exists a conic $\mathbf{{T}}^{0}$ that has double contact with $\mathbf{{S}}_{1},\mathbf{{S}}_{2},\mathbf{{S}}_{3}$.
@@
Here, the consistency condition of each cube face is that the four chords of contact from the four conics have a common point. This condition is automatically guaranteed if one picks a consistent branch of double contact conics with two given conics (out of 3 branches).
@@
This article provides a completion of one of Roger Penrose’s proof ideas for the eight-conic theorem. The idea emerged during his graduate studies at Cambridge that the eight-conic configuration is a planar slice of eight quadrics in ring contact in 3D. (Two quadrics have ring contact if they are tangent to each other along a conic.) It turns out that the following Penrose’s eight-quadric theorem is much more intuitive to see how and why it holds, as shown later in this article:
@@
If quadrics in 3D are assigned to seven of the vertices of a combinatorial cube such that quadrics connected by an edge are in ring contact, then (subject to a mild consistency condition on the cube faces) there exists an eighth quadric which completes the cube.
@@
The face consistency condition is that the four planes containing the four rings of the ring contacts must meet along a common axis. One advantage of the 3D approach is that this face consistency condition presents a much less serious obstacle to completing the cube than the 2D face consistency condition.
@@
The article is structured as follows. Section 2 provides the necessary background information on conics and quadrics. In Section 3, we present the main theorems: the eight-conic theorem and its 3D counterpart, the eight-quadric theorem. The proof of the eight-quadric theorem is given in Section 3.4, with the eight-conic theorem derived as a corollary through an extrusion process explained in Section 4. Section 5 illustrates examples of the eight-quadric configuration, showing how its 2D slices or projections yield classical theorems in the projective plane. The final sections of the article provide references to related works and future perspectives.
@@ 2.1 Conics and the projective plane
We recall that a point $P$ in the projective plane $\mathbb{R}P^{2}$ can be represented by an equivalence class of vectors in a 3-dimensional real [^7] We work primarily over the real field but allow complex coefficients when they arise through algebraic closure. vector space $\mathtt{V}$:
@@
Points are said to be “collinear” if they lie on the same line; lines are “copunctual” if they pass through the same point.
@@
for $v\in\mathtt{V}^{*}\setminus\{0\}$. We could equally well have started by describing lines according to Eq. (3) and proceeded to points by describing them as the sets of incident lines:
@@
thus, the roles of the two are completely interchangeable: the principle of duality. The symmetric relationship of the two can be expressed with the notation
@@
A point-line correlation is an invertible map $\mathbb{R}P^{2}\mapsto{\mathbb{R}P^{2}}^{*}$ induced by a linear transformation $\mathtt{V}\mapsto\mathtt{V}^{*}$. The fundamental theorem of projective geometry states that this is the most general type of mapping between the two, which always sends any three collinear points to three copunctual lines. A line-point correlation is a similarly defined map ${\mathbb{R}P^{2}}^{*}\mapsto{\mathbb{R}P^{2}}$.
@@
The linear transformation defining a point-line correlation can then be represented by a matrix $A$ relative to the bases $\mathcal{B}$ and $\mathcal{B}^{*}$, and vice versa for a line-point correlation. The matrix $cA$ for $c\neq 0$ clearly represents the same correlation (note, however, that changing the relative scaling of the vectors representing $P_{j}$ leads to a different matrix). A point-line correlation has two naturally associated line-point correlations: its inverse, which has matrix $A^{-1}$, and the mapping of a line $p$ to the intersection point of any two lines mapped to by two points lying on $p$, which is well-defined and has matrix $\mathrm{cof}A$, the cofactor matrix of $A$. By Cramer’s rule,
@@
where the superscript T denotes the transpose, therefore these two associated line-point correlations coincide exactly when the point-line correlation is represented by a symmetric matrix $A=A^{T}$ which is known as a polarity. A polarity can also be characterized as a correlation that is also an involution.
If $A$ is the matrix of a polarity and $P$ is a point, which we identify with the column vector of a representative in the basis $(P_{0},P_{1},P_{2})$, then the line $p$ polar to $P$ is represented by a column vector, which we again identify with $p$ itself, in the basis $(p_{0},p_{1},p_{2})$:
$p=AP$
@@
$p^{T}P=P^{T}AP=0$
The locus of such points is a regular point-wise conic which is uniquely defined by the polarity and vice versa, both having a unique representative symmetric matrix (up to scaling) in a given basis. The line to which a point of a point-wise conic is both polar and incident is its tangent line. Dual to this, the locus of lines incident to their polar points is called a regular line-wise conic and we refer to these points as the turning points of the line-wise conic. When convenient and no ambiguity can result, we sometimes identify the matrix representing a conic in a given basis with the conic itself.
Given a regular point-wise conic, there is a uniquely associated regular line-wise conic, the lines of the latter being the tangent lines of the points of the former. If the first conic is represented by the matrix $A$, the second is represented by the inverse matrix $A^{-1}$. A degenerate point-wise conic is a conic defined in the same way but by a non-invertible symmetric matrix. In this case, three collinear points will still map to three copunctual lines when all images are defined, however, for any point $P$ whose representatives lie in the kernel of the defining matrix $A$, the image of $P$ under the degenerate polarity is geometrically undefined, since the zero vector does not represent any line in ${\mathbb{R}P^{2}}^{*}$, moreover, if $Q$ is any other point, then
@@
  • If $\mathrm{ker}A$ has dimension $1$ then it constitutes one equivalence class in $\mathbb{R}P^{2}$ so there is a unique such point. In this case, the points of a degenerate point-wise conic constitute a line pair with $P$ as their meeting point, called the carrying point of the line pair. If these two lines are called $p$ and $q$ then the matrix is given by
@@
  • If $\mathrm{ker}A$ has dimension $2$ then it constitutes the representatives of the points on a line in the projective plane. In this case the points of a degenerate point-wise conic constitute a line, referred to for algebraic reasons as a double line. If this line is called $p$ then the matrix is given by
@@
Dual to this, degenerate line-wise conics can be defined:
  • If $\mathrm{ker}A$ has dimension $1$ then it constitutes one equivalence class in ${\mathbb{R}P^{2}}^{*}$ so there is a unique line $p$ represented by vectors in the kernel. In this case the lines of a degenerate line-wise conic constitute a point pair with $p$ as their joining line, called the carrying line of the point pair. If these two points are called $P$ and $Q$ then the matrix is given by
@@
  • If $\mathrm{ker}A$ has dimension $2$ then it constitutes the representatives of the lines through a point in the projective plane. In this case, the lines of a degenerate line-wise conic constitute a point, referred to for algebraic reasons as a double point. If this point is called $P$ then the matrix is given by
@@
2.1.1 Pencils of conics
The linear span of two conics $A$ and $B$ is called a pencil of conics. A pencil that contains a regular conic is called a regular pencil. We will need the following lemma in Sect. 3.2:
Lemma 2.1 (Regular pencil).
A regular pencil of conics contains at most 3 distinct degenerate conics.
Proof.
Assume $A$ is regular. The pencil, with the exception of $A$, can be written $A+tB,t\in\mathbb{R}$. $A+tB$ is a degenerate conic exactly when $|A+tB|=0$. The left-hand side is a cubic polynomial in $t$, hence has at most 3 roots. ∎